# Graph Topology Identification Based on Covariance Matching

### v1

@@ I Introduction
In addition to directed graphs, undirected graphs have also been studied within the SEM framework, and various methods have been proposed to address structure learning. For instance, approaches based on spectral templates (SpecTemp) leverage the eigen-decomposition of the data covariance matrix to estimate the underlying undirected topology [33]. These methods do not assume a specific graph filter or restrict its functional form, and thus often require a large number of samples. Another eigen-based method is the covariance matching (CovMatch) framework introduced in [17], which streamlines the estimation process by converting the graph estimation task into a binary variable inference problem. However, CovMatch formulated in [17] only enforces the hollowness constraint (no self-loops) and does not use additional priors such as sparsity, so accurate recovery typically requires very large sample sizes even SEM structure is explicitly considered. Building on this idea, in this work we extend CovMatch to more general settings, demonstrating that it effectively handles both undirected graphs and general directed graphs with only a sparsity constraint on the number of active edges. Our proposed approach easily integrates additional constraints when needed and, even without explicit constraints, can still recover diverse types of graphs, including DAGs and positively weighted directed graphs.

### v2

# A Covariance Matching Approach to Graph Topology Identification

@@ I Introduction
In addition to directed graphs, undirected graphs have also been studied extensively in GSP and various methods have been proposed to address structure learning. For instance, approaches based on spectral templates (SpecTemp) leverage the eigen-decomposition of the data covariance matrix to estimate the underlying undirected topology [33]. When the graph stationarity does not hold, a related line of work leverages richer second-order information—typically through multiple input–output diffusion pairs—to identify an unknown graph filter from covariance relations and then infer a sparse graph shift operator (GSO) via dedicated optimization, which has been developed for undirected graphs [35] and directed graphs via optimization on the Stiefel manifold [34]. These methods do not assume a specific graph filter or restrict its functional form, and thus often require a large number of samples. Another covariance-based method is the covariance matching (CovMatch) framework, which was recently introduced in [17]. It streamlines the estimation process by converting the graph estimation task into a binary variable inference problem. However, this framework only enforces the hollowness constraint (i.e., no self-loops) and does not use additional priors such as sparsity. Hence, accurate recovery typically requires very large sample sizes, even if the SEM structure is explicitly considered. In this work, we build on [17] and generalize CovMatch to handle both undirected graphs and general directed graphs, with only a sparsity constraint on the number of active edges. Our proposed approach easily integrates additional constraints when needed and, even without explicit constraints (except for sparsity), can still recover diverse types of graphs, including DAGs and positively weighted directed graphs.